# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ A.2 Factor analysis
Table 6 summarizes the different configurations for FA. We assume a Gaussian prior on the mean components, and an inverse gamma prior on the elements of the diagonal covariance matrix $\bm{\Psi}$. For the factor loading matrix $\bm{W}$, independent normal and Laplace priors are investigated. Furthermore, we use a normal prior on the latent factors ${\bm{z}}^{(i)}$ in five cases and a Laplace prior in one case. We vary the number of samples $K$ per dataset ${\bm{x}}$, the dimensionality $P$ of each data point, as well as the dimensionality $\bm{z}_{dim}$.
1
2Initialize $\mathcal{D}\leftarrow\o$;
3 for $i=1,\ldots N$ do
4       draw $\bm{\mu}_{i}\sim P^{\bm{\mu}}$ ;
5       draw $\bm{\Psi}_{i}\sim P^{\bm{\Psi}}$ ;
6       draw $\bm{W}_{i}\sim P^{\bm{W}}$ ;
7       draw ${\bm{z}}^{(i)}\sim P^{{\bm{z}}}$ ;
8       for $j=1,\ldots,K$ do
9             draw $\bm{x}_{i,j}\sim\mathcal{N}(\bm{W}_{i}{\bm{z}}^{(i)}+\bm{\mu}_{i},\bm{\Psi}_{i})$;
10
11       end for
12      $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}$;
13
14 end for
15
Algorithm 3 Generation of synthetic data for FA
@@ 12      $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}$;
Scenario 1 $50$ $3$ $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,1)$ $3$
Scenario 2 $50$ $3$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,10)$ $\mathcal{N}(0,1)$ $3$
Scenario 3 $25$ $5$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\mathcal{N}(0,1)$ $3$
Scenario 4 $25$ $15$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\mathcal{N}(0,1)$ $5$
Scenario 5 $25$ $5$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\text{Laplace}(0,3)$ $\mathcal{N}(0,1)$ $3$
Scenario 6 $25$ $5$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\text{Laplace}(0,1)$ $3$
@@ A.3 Gaussian mixture models
1
2Initialize $\mathcal{D}\leftarrow\o$;
3 for $i=1,\ldots N$ do
4       draw $\bm{\phi}_{i}\sim P^{\bm{\phi}}$ ;
5
6      for $m=1,\ldots,M$ do
7             for $l=1,\ldots,L$ do
8                   draw $\sigma^{2}_{i,m,l}\sim P^{\bm{\sigma}^{2}}$ ;
9                   draw $\mu_{i,m,l}\sim P^{\bm{\mu}|\bm{\sigma}^{2}_{i,m,l}}$ ;
10
11             end for
12
13       end for
14
15      for $j=1,\ldots,K$ do
16             draw $\omega_{i,j}\sim\text{Cat}(\bm{\phi}_{i})$;
17
18            draw $\bm{x}_{i,j}\sim\mathcal{N}(\bm{\mu}_{i,\bm{\omega}_{i,j}},\bm{\sigma}^{2}_{i,\bm{\omega}_{i,j}})$;
19
20       end for
21      set ${\bm{z}}^{(i)}\vcentcolon=\left((\sigma^{2}_{i,m,l},\mu_{i,m,l})\right)_{m,l=1}^{M,L}$;
22       $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}$;
23
24 end for
25
Algorithm 4 Generation of synthetic data for a GMM.

### v2

@@ A.2 Factor Analysis
Table 6 summarizes the different configurations for FA. We assume a Gaussian prior on the mean components, and an inverse gamma prior on the elements of the diagonal covariance matrix $\bm{\Psi}$. For the factor loading matrix $\bm{W}$, independent normal and Laplace priors are investigated. Furthermore, we use a normal prior on the latent factors $\bm{z}_{i}$ in five cases and a Laplace prior in one case. We vary the number of samples $K$ per dataset $\bm{x}$, the dimensionality $P$ of each data point, as well as the dimensionality $\bm{z}_{dim}$.
@@
Scenario 1 $50$ 33 $\mathcal{N}(0,1)$ $\text{IG}(5,1)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,1)$ 33
Scenario 2 $50$ 33 $\mathcal{N}(0,0.1)$ $\text{IG}(5,1)$ $\text{Laplace}(0,10)$ $\mathcal{N}(0,1)$ 33
Scenario 3 $25$ 55 $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\mathcal{N}(0,1)$ 33
Scenario 4 $25$ $15$ $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\mathcal{N}(0,1)$ 55
Scenario 5 $25$ 55 $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\text{Laplace}(0,3)$ $\mathcal{N}(0,1)$ 33
Scenario 6 $25$ 55 $\mathcal{N}(0,0.1)$ $\text{IG}(5,2)$ $\mathcal{N}(0,3)$ $\text{Laplace}(0,1)$ 33
Algorithm 3 Generation of synthetic data for FA
0: Number of datasets $N$, number of samples $K$, and distributions $P^{\bm{\mu}},P^{\bm{\Psi}},P^{\bm{W}},P^{\bm{z}}$.
0: A dataset $\mathcal{D}$ containing $(\bm{x}_{i},\bm{z}_{i})$ for $i=1,\dots,N$.
1: Initialize $\mathcal{D}\leftarrow\varnothing$
2:for $i=1\to N$ do
3:  Draw $\bm{\mu}_{i}\sim P^{\bm{\mu}}$
4:  Draw $\bm{\Psi}_{i}\sim P^{\bm{\Psi}}$
5:  Draw $\bm{W}_{i}\sim P^{\bm{W}}$
6:  Draw $\bm{z}_{i}\sim P^{\bm{z}}$
7:  for $j=1\to K$ do
8:   Draw $\bm{x}_{i,j}\sim\mathcal{N}\!\bigl{(}\bm{W}_{i}\,\bm{z}_{i}+\bm{\mu}_{i},\;\bm{\Psi}_{i}\bigr{)}$
9:  end for
10:  Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{\bigl{(}\bm{x}_{i},\bm{z}_{i}\bigr{)}\}$
11:end for